OpenStudy (anonymous):

What is the exact value for the expression sqrt(171)- sqrt(19)+sqrt(76)

OpenStudy (anonymous):

Check if the values under the square root have a least common factor.

OpenStudy (anonymous):

no nothing goes into 19

OpenStudy (anonymous):

Remember that 19 is a factor of 19

OpenStudy (anonymous):

so would it just be 19

OpenStudy (anonymous):

Ok then split the other larger values up for exampe sqrt(50)=sqrt(25*2)

OpenStudy (anonymous):

sqrt(171)= (19*9) srt(19)= (19*1) sqrt(76)=( 19*4)

OpenStudy (anonymous):

Now sqrt(a*b)=sqrt(a)*sqrt(b)

OpenStudy (anonymous):

19 sqrt(9) 19 sqrt(1) 19 sqrt(4)

OpenStudy (anonymous):

No you still need the sqrt over your 19s

OpenStudy (anonymous):

o ok sqrt (19) sqrt(9) sqrt(19) sqrt(1) sqrt(19) sqrt(4)

OpenStudy (anonymous):

Now some of those can be simplified. Then substitute them back into your original equation

OpenStudy (anonymous):

wouldn't be 17.43558577416269

OpenStudy (anonymous):

That's not an exact value. When values that are not perfect squares are under square roots, you cant simplify any more without approximating

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

oh my goodness i would have nvr thought to do this ans is sqrt 304

OpenStudy (anonymous):

would you get 4 sqrt(19)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you